(e) A solid cube of mass 100 g and density 2.70 x 103 kg m-3 is subjected to a shear force of 5.00 kN. The shear deformation is 5.00 μm. Find the shear moduļus.
Answer #1

Shear Modulus, G =FL/Ax

Volume of cube, V = L​​​​​​3​​​ = mass/density

Therefore,

L = (0.100/2700)1/3 = 3.33 × 10-2m

Area, A = L​​​​​​2​​​

Therefore,

G = F/Lx

G = (5000)/(3.33 × 10-2m)(5.00 ×10-6m)

G = 3×1010 N/m​​​​​​2